# Estimating Fold Changes from Partially Observed Outcomes with Applications in Microbial Metagenomics

### v1

@@ Abstract
We consider the problem of estimating fold-changes in the expected value of a multivariate outcome that is observed subject to unknown sample-specific and category-specific perturbations. We are motivated by high-throughput sequencing studies of the abundance of microbial taxa, in which microbes are systematically over- and under-detected relative to their true abundances. Our log-linear model admits a partially identifiable estimand, and we establish full identifiability by imposing interpretable parameter constraints. To reduce bias and guarantee the existence of parameter estimates in the presence of sparse observations, we apply an asymptotically negligible and constraint-invariant penalty to our estimating function. We develop a fast coordinate descent algorithm for estimation, and an augmented Lagrangian algorithm for estimation under null hypotheses. We construct a model-robust score test, and demonstrate valid inference even for small sample sizes and violated distributional assumptions. The flexibility of the approach and comparisons to related methods are illustrated via a meta-analysis of microbial associations with colorectal cancer.
@@ 1 Introduction
Communities of microorganisms known as microbiomes have become a focus of study across multiple scientific disciplines, including in many areas of human health and disease (Young, 2017; Yamashita and Takeshita, 2017; Shreiner et al., 2015; Moffatt and Cookson, 2017). Modern microbiome research leans heavily on high-throughput sequencing techniques, which enable researchers to identify large numbers of microbial taxa on the basis of genetic material detected in a given environment. In this paper, we consider the problem of making statistical statements about changes in the “absolute” concentrations of microbes in an environment using only the “relative” information provided by marker gene and whole genome shotgun sequencing. Our approach also accounts for the observation that different microbes are consistently over- or under-counted (McLaren et al., 2019).
Most generally, this paper introduces a method for inference on ratios of means of a nonnegative multivariate outcome observed only up to unknown sample-specific and category-specific scaling terms. To our knowledge no similar methods currently exist, but a related niche in microbiome science is occupied by a range of regression models that rely on log-ratio transformations to target a similar estimand. These include ANCOM and its variants, which attempt to distinguish sampling, structural, and “outlier” zero values in high-throughput sequencing data prior to transformation and modeling (Mandal et al., 2015; Kaul et al., 2017; Lin and Peddada, 2020); and ALDEx2, which employs a Bayesian approach to impute nonzero values before transformation (Fernandes et al., 2014).
@@
ANCOM and ALDEx2 fall into the larger category of “differential abundance” methods, which include LEfSe, DESeq2, edgeR, metagenomeSeq, IFAA, and corncob (Segata et al., 2011; Love et al., 2014; Robinson et al., 2010; Paulson et al., 2013; Li et al., 2021; Martin et al., 2020), as well as many others (see Nearing et al. (2022) for a recent empirical comparison). DESeq2 and edgeR were originally developed for RNAseq data and attempt to address excess-Poisson dispersion via a negative binomial likelihood (Love et al., 2014; Robinson et al., 2010). By contrast, metagenomeSeq was developed specifically for microbiome analysis and employs a zero-inflated Gaussian model motivated by the high degree of sparsity in microbiome datasets (Paulson et al., 2013). LEfSe, also developed for microbiome analysis, attempts to identify “metagenomic biomarkers” via a Kruskal-Wallis test followed by Wilcoxon tests on normalized sequencing data (Segata et al., 2011), and IFAA fits a log-linear zero-inflated Gaussian model by identifying a reference taxon whose abundance is conditionally independent of covariates (Li et al., 2021). Corncob (Martin et al., 2020) models both taxon relative abundances and dispersion of relative abundances via a beta-binomial model with logit link functions. We note that although the regression models above are all referred to as differential abundance methods (Nearing et al., 2022), they do not target a common estimand and hence do not produce strictly comparable inference. Furthermore, most of these methods target estimands defined in terms of microbial proportions, while our approach targets an estimand defined on the total microbial abundance scale (e.g., taxon-specific cell counts, DNA concentration, or copy number).
@@
Our manuscript is structured as follows: we introduce our model and establish identifiability of equivalence classes of parameters in Section 2. We discuss estimation via a generalized estimating equations framework in Section 3, and inference via model misspecification-robust testing in Section 4. We present simulation results pertaining to Type 1 and 2 error rates in Section 5 and consider a data analysis in Section 6. We conclude with a discussion in Section 7. Open-source software implementing the method is available at github.com/statdivlab/radEmu, and code to reproduce the analyses are available at github.com/statdivlab/radEmu_supplementary.
@@ 2.1 Motivation
Our work is motivated by settings in which an “ideal” dataset $W_{ij}\geq 0$ (samples $i=1,\dots,n$ and categories $j=1,\dots,J$) is not observed, but a distorted proxy $Y_{ij}\geq 0$ is available. Observing $Y$ rather than $W$ represents a loss of information in two critical ways. Firstly, the magnitude of $Y_{i}\in\mathbb{R}^{J}_{\geq 0}$ is uninformative with respect to the magnitude of $W_{i}\in\mathbb{R}^{J}_{\geq 0}$. In addition, the composition of $Y_{i}$ may be systematically distorted relative to $W_{i}$ as a result of differential detection. That is, observations from category $j_{1}$ may be more readily observed than observations from category $j_{2}$, and the magnitude of the detection effects are unknown. Our goal is to estimate fold-differences in $\mathbb{E}W_{\cdot j}$ across covariates, that is, $\mathbb{E}W_{\{i:X_{i}=x\}j}/\mathbb{E}W_{\{i:X_{i}=x^{\prime}\}j}$. We formalize our model and approach to estimating fold-differences in Section 2.2.
We our motivating example of microbial communities, $W_{ij}$ represents the (unobserved) cell or DNA concentration of microbial taxon $j$ in sample $i$, which are prohibitively time- and labor-intensive to observe for all taxa $j=1,\ldots,J$ (Williamson et al., 2022). However, it is relatively straightforward to perform high-throughput sequencing to produce data $Y_{ij}$, such as the (observed) read depth of taxon $j$ in sample $i$ in shotgun sequencing, or the number of times sequence variant $j$ is observed in sample $i$ based on amplicon sequencing. In the case-control microbiome studies considered in Section 6, we are interested in estimating the fold-difference in the mean cell concentration of taxa $j$ when comparing cases to otherwise similar controls.
@@ 2.2 Model statement and parameter identifiability
Suppose that for pairs $(X_{i},W_{i})$ with $X_{i}\in\mathbb{R}^{p}$ and $W_{i}\in\mathbb{R}^{J}_{\geq 0}$,
$\log\mathbb{E}[W_{i}|X_{i},\beta^{\star}]=X_{i}\beta^{\star}.$ (1)
Our target of inference is $\beta^{\star}\in\mathbbm{R}^{p\times J}$. Unfortunately, however, we do not observe $W_{i}$, and so we cannot directly estimate $\beta^{\star}$ via (1). Instead, we instead observe a perturbed outcome $Y_{i}\in\mathbb{R}^{J}_{\geq 0}$ such that
$\log\mathbb{E}[Y_{i}|X_{i},\beta^{\star},z_{i},\delta]=z_{i}\mathbf{1}_{J}^{T}+X_{i}\beta^{\star}+\delta^{T},$ (2)
where sample-specific scalings $z_{i}\in\mathbb{R}$ and category-specific detection effects $\delta\in\mathbbm{R}^{J}$ are unknown. Hence, the mean of $Y_{ij}$ is equal to the mean of $W_{ij}$ subject to scaling by the sample-specific factor $\exp(z_{i})$ and category-specific factor $\exp(\delta_{j})$.
As our target of inference is $\beta^{\star},$ we consider $\mathbf{z}=[z_{1},\dots,z_{n}]^{T}$ and $\delta$ as nuisance parameters. $\delta$ is straightforward to address if $\mathbf{1}_{n}$ lies in the column space of $\mathbf{X}=[X_{1}^{T},\dots,X_{n}^{T}]^{T}$, as is typically the case in practice. Without loss of generality, we define $\mathbf{X}=[\mathbf{1}_{n}~{}\mathbf{D}]\in\mathbb{R}^{n\times p}$ and $\mathbf{Y}=[Y_{1}^{T},\dots,Y_{n}^{T}]^{T}\in\mathbb{R}^{n\times J}$, and rewrite model (14) as follows:
$\log\mathbb{E}[\mathbf{Y}|\mathbf{X}]$ $=\mathbf{z}\mathbf{1}_{J}^{T}+\mathbf{1}_{n}\beta_{0}+\mathbf{D}\beta^{\star(-0)}=\mathbf{z}\mathbf{1}_{J}^{T}+\mathbf{X}\beta$ (3)
for $\beta_{0}=\beta_{0}^{\star}+\delta^{T}$ and $\beta=[\beta_{0}^{T},\beta_{1}^{\star T},\dots,\beta_{p-1}^{\star T}]^{T}$, for $\beta_{k}^{\star}\in\mathbb{R}^{1\times J}$ the $k+1$-th row of $\beta^{\star}$. In particular, because $\beta_{0}$ depends on the nuisance detection effects $\delta$, it has no scientifically meaningful interpretation, but $\beta_{1},\ldots,\beta_{p-1}$ contain useful information regarding $\beta_{1}^{\star},\ldots,\beta_{p-1}^{\star}$, as discussed in more detail below.
@@
$z_{i}\mathbf{1}_{J}^{T}+X_{i}(\beta+\alpha\mathbf{1}_{J}^{T})$ $=z^{\dagger}_{i}\mathbf{1}_{J}^{T}+X_{i}\beta$ (4)
for $z^{\dagger}_{i}:=(z_{i}+X_{i}\alpha)$. Hence if $\beta^{\dagger}=\beta+\alpha\mathbf{1}_{J}^{T}$ for some $\alpha\in\mathbb{R}^{p}$, we can always find a $z_{i}^{\dagger}\in\mathbb{R}$ such that
$\log\mathbb{E}[Y_{i}|X_{i},\beta,z_{i}]=\log\mathbb{E}[Y_{i}|X_{i},\beta^{\dagger},z_{i}^{\dagger}],$ (5)
and therefore mean function (14) is not identifiable in $\beta$ and $z_{i}$. However, mean function (14) is partially identifiable because we can define equivalence classes of $\beta$’s such that for $\beta$ and $\beta^{\dagger}$ in the same equivalence class, there always exist $z_{i}$ and $z_{i}^{\dagger}$ such that (5) holds, and for $\beta$ and $\beta^{\dagger}$ in distinct classes, there never exist such $z_{i}$ and $z_{i}^{\dagger}$. We now make the form of these classes explicit and show that they have the properties discussed above.
@@
$G_{b}=\{\beta\in\mathbbm{R}^{p\times J}:\beta=b+\alpha\mathbf{1}_{J}^{T}\text{for some }\alpha\in\mathbb{R}^{p}\}.$
@@
If $\mathbf{X}$ has full column rank, for any $\beta,\beta^{\dagger}\in\mathbb{R}^{p\times J}$ there exist $\mathbf{z}\in\mathbb{R}^{n}$ and $\mathbf{z}^{\dagger}\in\mathbb{R}^{n}$ such that under model (14), $\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta,\mathbf{z}]=\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta^{\dagger},\mathbf{z}^{\dagger}]$ if and only if $\beta^{\dagger}\in G_{\beta}$.
@@
We prove Lemma 1 in Supporting Information Section 1. Lemma 1 demonstrates that while we cannot uniquely estimate $\beta$ from data generated according to (14), we can estimate it up to the addition of constant terms to each of its rows. That is, while we cannot estimate differences across levels of covariate $X_{\cdot k}$ in the log-means of $\mathbf{W}$ given only data $\mathbf{Y}$, we can estimate relative differences in the log-mean of $\mathbf{W}$ in category $j_{1}$ compared to in category $j_{2}$. In microbiome sequencing studies, this represents a small loss of information because researchers typically aim to identify microbial taxa that show markedly different abundances across covariate groups of interest when compared to the typical magnitude of these differences. Therefore, estimating the equivalence class of $\beta$ answers important scientific questions about microbial abundances.
@@
We estimate the equivalence class of $\beta$ by imposing identifiability constraints of the form $g(\beta_{k})=0$ on the rows $\beta_{k}$ of $\beta$ such that each $G_{\beta}$ contains only a single unique element satisfying these constraints. Fitted means produced by our model are invariant to choice of constraint, but this choice does change the interpretation of elements of $\beta$. For example, under constraint $\beta_{kj}=0$, $\beta_{kj^{\prime}}$ represents $\log\mathbb{E}[W_{\cdot j^{\prime}}|X_{\cdot}=\mathbf{x}_{x_{k}+1}]-\log\mathbb{E}[W_{\cdot j^{\prime}}|X_{\cdot}=\mathbf{x}]$ minus the corresponding quantity for taxon $j$, where $\mathbf{x}_{x_{k}+1}$ is equal to $\mathbf{x}$ in all elements excepting the $k$th, which is increased by 1 unit. We equivalently interpret $e^{\beta_{kj^{\prime}}}$ as a fold-difference (multiplicative difference).
Figure 1: Because the mean function is only partially identifiable, true effect sizes $\beta$ cannot be estimated on an “absolute” scale (left). Full identifiability is established via a constraint function, allowing us to estimate the log-fold differences in true abundances across groups relative to typical differences. The above illustrates the favorable qualities of centering via the pseudo-Huber smoothed median (right) compared to centering by the mean (middle). $\beta_{j}$ for $j=2,3,4,5$ are unlikely to be of scientific interest (relative to $j=1,6$), and the smoothed median constrains these “uninteresting” taxa closer to zero than mean centering.
@@
If we aim to compare log-mean differences across levels of a covariate to typical such differences across taxa, we might choose to impose identifiability by requiring $\text{median}_{j=1,\ldots,J}(\beta_{kj})=0$. The robustness of the median is attractive: if few taxa exhibit large differences in mean concentration across levels of a covariate, this will in general minimally impact a median effect if most effects are small. By contrast, centering by a non-robust measure of location (e.g., mean) may result in unwanted behavior, because a single large effect $\beta_{kj}$ will pull the mean away from the bulk of taxa (Figure 1). However, the median is not a differentiable function of $\beta_{k}$ (its total derivative is not guaranteed to exist), which substantially complicates inference (see Section 4). Accordingly, we propose a smooth approximation to the median that shares its robustness properties, that is, the pseudo-Huber loss $g_{p}(\beta_{k})=\arg\min_{c}\sum_{j=1}^{J}h_{\delta}\left(\beta_{kj}-c\right)$ for $h_{\delta}(x)=\delta^{2}\left(\sqrt{1+\left(x/\delta\right)^{2}}-1\right)$ for prespecified parameter $\delta>0$ (Barron, 2019). The pseudo-Huber loss is approximately quadratic in a neighborhood of 0 and asymptotically approaches the function $|x|/\delta$ for large $x$. The parameter $\delta$ controls the size of the quadratic region around zero: as $\delta\downarrow 0$, $g_{p}(\beta_{k})$ approaches the median, and as $\delta\rightarrow\infty$, it approaches the mean. We propose $\delta=0.1$ as a reasonable default for many applications, though empiricially we find that care is warranted when covariates are on very different scales. While we propose $g_{p}$ as the default constrained function, any smooth $g$ satisfying $g(\beta_{k}+\alpha)=g(\beta_{k})+\alpha$ is reasonable, and the most appropriate constraint will depend on the scientific setting.

### v2

@@ Abstract
We consider the problem of estimating fold-changes in the expected value of a multivariate outcome observed with unknown sample-specific and category-specific perturbations. This challenge arises in high-throughput sequencing studies of the abundance of microbial taxa because microbes are systematically over- and under-detected relative to their true abundances. Our model admits a partially identifiable estimand, and we establish full identifiability by imposing interpretable parameter constraints. To reduce bias and guarantee the existence of estimators in the presence of sparse observations, we apply an asymptotically negligible and constraint-invariant penalty to our estimating function. We develop a fast coordinate descent algorithm for estimation, and an augmented Lagrangian algorithm for estimation under null hypotheses. We construct a model-robust score test and demonstrate valid inference even for small sample sizes and violated distributional assumptions. The flexibility of the approach and comparisons to related methods are illustrated through a meta-analysis of microbial associations with colorectal cancer.
@@ 1 Introduction
Communities of microorganisms (microbiomes) are studied in many scientific disciplines, including in multiple areas of human health and disease (Young, 2017; Yamashita and Takeshita, 2017; Shreiner et al., 2015; Moffatt and Cookson, 2017). Modern microbiome research leans heavily on high-throughput sequencing techniques, which enable researchers to identify large numbers of microbial taxa on the basis of genetic material detected in a given environment. In this paper, we consider the problem of making statistical statements about changes in the “absolute” concentrations of microbes in an environment using only the “relative” information provided by marker gene and whole genome shotgun sequencing. Our approach also accounts for the observation that different microbes are consistently over- or under-counted (McLaren et al., 2019).
@@
We introduce a method for estimation and inference on ratios of means of a nonnegative multivariate outcome observed only up to unknown sample-specific and category-specific perturbations (Figure 1). To our knowledge, this problem has not been studied in the statistical literature. Unknown perturbations are addressed by microbiome data analysts via a range of regression models that typically rely on log-based transformations. These include ANCOM-BC2, which delineates zero-valued outcomes into “structural” and non-structural zeroes prior to transformation and modeling (Lin and Peddada, 2024); and ALDEx2, which employs a Bayesian approach to impute nonzero values before transformation (Fernandes et al., 2014). ANCOM-BC2 and ALDEx2 fall into the larger category of “differential abundance” methods, which include DESeq2, edgeR, metagenomeSeq, IFAA, and corncob, as well as many others (see Nearing et al. (2022) for an empirical comparison). DESeq2 and edgeR were originally developed for RNAseq data and address excess-Poisson dispersion via a negative binomial likelihood (Love et al., 2014; Robinson et al., 2010). By contrast, metagenomeSeq was developed specifically for microbiome analysis and employs a zero-inflated Gaussian model motivated by the high degree of sparsity in microbiome datasets (Paulson et al., 2013). IFAA fits a log-linear zero-inflated Gaussian model by identifying a reference taxon whose abundance is conditionally independent of covariates (Li et al., 2021), and corncob (Martin et al., 2020) models both taxon relative abundances and dispersion of relative abundances via a beta-binomial model. Although the regression models above are all considered to be differential abundance methods (Nearing et al., 2022), they do not target a common estimand and hence do not produce strictly comparable inference.
@@
Our manuscript is structured as follows: we introduce our model and establish identifiability of equivalence classes of parameters in Section 2. We discuss estimation via a generalized estimating equations framework in Section 3, desirable asymptotic properties of our estimator in Section 4, and inference via model misspecification-robust testing in Section 5. We present simulation results pertaining to Type 1 and 2 error rates in Section 6 and consider a data analysis in Section 7. We conclude with a discussion in Section 8. Open-source software implementing the method is available at github.com/statdivlab/radEmu, and additional details and proofs are available in Supporting Information (SI).
Figure 1: Motivated by differential abundance in microbiome studies, we consider the general problem of estimating fold-differences in the mean of a multivariate outcome in the presence of measurement error. Specifically, when the expected value of the true outcome $W_{ij}$ is multiplicatively distorted by unknown sample-specific ($e^{z_{i}}$) and category-specific ($e^{\delta_{j}}$) terms, we show that $e^{\beta_{kj}}=\mathbb{E}W_{ij}(\mathbf{x}_{x_{k}+1})/\mathbb{E}W_{ij}(\mathbf{x})$ is not identifiable. However, we show that $\beta_{kj}-g(\beta_{k\cdot})$ is identifiable for appropriate choices of $g:\mathbb{R}^{J}\rightarrow\mathbb{R}$, a centering function. Reasonable choices include mean-centering, reference group-centering, and smoothed median-centering. The choice of centering function impacts the interpretation of the resulting estimates, but our proposed estimation procedure is $g$-invariant.
@@ 2.1 Motivation
Our work is motivated by settings in which an “ideal” dataset $W_{ij}\geq 0$ (samples $i=1,\dots,n$ and categories $j=1,\dots,J$) is not observed, but a distorted proxy $Y_{ij}\geq 0$ is available. Observing $Y$ rather than $W$ represents a loss of information in two critical ways (Figure 1). Firstly, the magnitude of $Y_{i}\in\mathbb{R}^{J}_{\geq 0}$ is uninformative with respect to the magnitude of $W_{i}\in\mathbb{R}^{J}_{\geq 0}$. In addition, the composition of $Y_{i}$ may be systematically distorted relative to $W_{i}$ as a result of differential detection. That is, observations from category $j_{1}$ may be more readily observed than observations from category $j_{2}$, and the magnitude of the detection effects are unknown. Our goal is study the estimability of fold-differences in $\mathbb{E}W_{\cdot j}$ across covariates, that is, $\mathbb{E}W_{\{i:X_{i}=x\}j}/\mathbb{E}W_{\{i:X_{i}=x^{\prime}\}j}$.
In our motivating example of microbial communities, $W_{ij}$ represents the (unobserved) cell or DNA concentration of microbial taxon $j$ in sample $i$, which are prohibitively time- and labor-intensive to observe for all taxa $j=1,\ldots,J$ (Williamson et al., 2022). However, it is relatively straightforward to perform high-throughput sequencing to produce data $Y_{ij}$, such as the (observed) read coverage of taxon $j$ in sample $i$ in shotgun sequencing, or the number of times sequence variant $j$ is observed in sample $i$ based on amplicon sequencing. In the microbiome meta-analysis considered in Section 7, we are interested in estimating the fold-difference in the mean cell concentration of taxon $j$ when comparing cancer-diagnosed cases to otherwise similar non-cancer controls, using shotgun sequencing data.
@@ 2.2 Model statement and parameter identifiability
Suppose that
$\log\mathbb{E}[\mathbf{W}|\mathbf{X}]=\mathbf{X}\beta.$ (1)
for $\mathbf{W}\in\mathbb{R}^{n\times J}_{\geq 0}$ and $\mathbf{X}\in\mathbb{R}^{n\times p}$. Our target of inference is $\beta\in\mathbbm{R}^{p\times J}$. Unfortunately, however, we do not observe $\mathbf{W}$, and so we cannot directly estimate $\beta$ via (1). Instead, we observe a perturbed outcome $\mathbf{Y}\in\mathbb{R}^{n\times J}_{\geq 0}$ such that
$\log\mathbb{E}[\mathbf{Y}|\mathbf{X}]=\mathbf{z}\mathbf{1}_{J}^{T}+\mathbf{1}_{n}\delta^{T}+\mathbf{X}\beta$ (2)
where sample-specific scalings $\mathbf{z}\in\mathbb{R}^{n\times 1}$ and category-specific detection effects $\delta\in\mathbb{R}^{J\times 1}$ are unknown (see Figure 1). Hence, the mean of $Y_{ij}$ is equal to the mean of $W_{ij}$ subject to scaling by the sample-specific factor $\exp(z_{i})$ and category-specific factor $\exp(\delta_{j})$.
As our target of inference is $\beta,$ we consider $\mathbf{z}$ and $\delta$ as nuisance parameters. $\delta$ is straightforward to address if $\mathbf{1}_{n}$ lies in the column space of $\mathbf{X}$, as is typically the case in practice. Without loss of generality, we consider $\mathbf{X}$ to have the form $[\mathbf{1}_{n}~{}\mathbf{D}]$, and rewrite model (2) as follows:
$\log\mathbb{E}[\mathbf{Y}|\mathbf{X}]=\mathbf{z}\mathbf{1}_{J}^{T}+\mathbf{X}\beta^{\star}$ (3)
for $\beta^{\star}=[\beta_{0}+\delta,\beta_{1},\dots,\beta_{p-1}]^{T}$, for $\beta_{k}\in\mathbb{R}^{J\times 1}$ the $k+1$-th row of $\beta$. In particular, because $\beta_{0}^{\star}$ depends on the nuisance detection effects $\delta$, it has no scientifically meaningful interpretation. We show below that while $\beta_{kj}$ for $k\geq 1$ is not identifiable from data $\mathbf{Y}$, $\beta_{kj}-g(\beta_{k})$ is identifiable for appropriate centering functions $g$.
@@
$\mathbf{z}\mathbf{1}_{J}^{T}+\mathbf{X}(\beta^{\star}+\alpha\mathbf{1}_{J}^{T})$ $=\mathbf{z}^{\dagger}\mathbf{1}_{J}^{T}+\mathbf{X}\beta^{\star}$ (4)
for $\mathbf{z}^{\dagger}:=(\mathbf{z}+\mathbf{X}\alpha)$. Hence if $\beta^{\dagger}=\beta^{\star}+\alpha\mathbf{1}_{J}^{T}$ for some $\alpha\in\mathbb{R}^{p}$, we can always find a $z_{i}^{\dagger}\in\mathbb{R}$ such that
$\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta^{\star},\mathbf{z}]=\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta^{\dagger},\mathbf{z}^{\dagger}],$ (5)
and therefore mean function (3) is not identifiable in $\beta^{\star}$ and $\mathbf{z}$. However, mean function (3) is partially identifiable because we can define equivalence classes such that for $\beta^{\star}$ and $\beta^{\dagger}$ in the same equivalence class, there always exist $\mathbf{z}$ and $\mathbf{z}^{\dagger}$ such that (5) holds, and for $\beta^{\star}$ and $\beta^{\dagger}$ in distinct classes, there never exist such $\mathbf{z}$ and $\mathbf{z}^{\dagger}$. We prove the following result in Supporting Information (Section 9).
@@
$G_{b}=\{\beta^{\star}\in\mathbbm{R}^{p\times J}:\beta^{\star}=b+\alpha\mathbf{1}_{J}^{T}\text{ for some }\alpha\in\mathbb{R}^{p}\}.$
@@
If $\mathbf{X}$ has full column rank, for any $\beta^{\star},\beta^{\dagger}\in\mathbb{R}^{p\times J}$ there exist $\mathbf{z}\in\mathbb{R}^{n}$ and $\mathbf{z}^{\dagger}\in\mathbb{R}^{n}$ such that under model (3), $\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta^{\star},\mathbf{z}]=\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta^{\dagger},\mathbf{z}^{\dagger}]$ if and only if $\beta^{\dagger}\in G_{\beta}$.
@@
Lemma 1 demonstrates that while we cannot uniquely estimate $\beta_{1},\ldots,\beta_{p-1}$ from data generated according to (2), we can estimate them up to the addition of constant terms. That is, while we cannot estimate differences across levels of covariate $X_{\cdot k}$ in the log-means of $W_{\cdot j}$ given only data $\mathbf{Y}$, we can estimate relative differences in the log-mean of $W_{\cdot j_{1}}$ compared to $W_{\cdot j_{2}}$. In genomics studies, this represents a small loss of information because researchers typically aim to identify categories that show markedly different abundances across covariate groups of interest. That is, even if all categories differ in abundance across covariate groups, the categories of most interest are those with the largest differences. Therefore, even if we cannot identify $\beta_{1},\ldots,\beta_{p-1}$ uniquely, estimating their equivalence classes can answer questions about differences in unobserved true abundances $\mathbf{W}$ across groups defined by $\mathbf{X}$.
@@
We estimate the equivalence class of $\beta_{1},\ldots,\beta_{p-1}$ by imposing identifiability constraints of the form $g(\beta_{k})=0$ such that each $G_{\beta}$ contains only a single unique element satisfying these constraints. As is typical of overparameterized models (e.g., one-way ANOVA in the absence of constraints on category means $\mu+\alpha_{i}$), the choice of constraint changes the interpretation of the parameters. For example, under constraint $\beta_{kj^{\prime}}=0$, $\beta_{kj}$ represents $\log\mathbb{E}[W_{\cdot j}|X_{\cdot}=\mathbf{x}_{x_{k}+1}]-\log\mathbb{E}[W_{\cdot j}|X_{\cdot}=\mathbf{x}]$ minus the corresponding quantity for taxon $j^{\prime}$, where $\mathbf{x}_{x_{k}+1}$ is equal to $\mathbf{x}$ in all elements excepting the $k$th, which is increased by 1 unit. We equivalently interpret $e^{\beta_{kj}}$ as a fold-difference (multiplicative difference).
Example 1: We contrast two scenarios: assuming that $\beta_{kj^{\prime}}=0$, and imposing the constraint that $\beta_{kj^{\prime}}=0$. In situations where a “reference category” is known to have $\beta_{kj^{\prime}}=0$, the constraint $\beta_{kj^{\prime}}=0$ means that $\beta_{kj}$ can be interpreted as the log-fold difference in $\mathbb{E}W_{\cdot j}$ across groups that differ by one unit in $X_{\cdot k}$. Therefore, knowledge of a reference category allows interpretation of parameters as fold-differences in mean abundances $W$. In contrast, when no such category is known, $\beta_{kj}$ is not identifiable from data $\mathbf{Y}$. Imposing the constraint $\beta_{kj^{\prime}}=0$ means that $\beta_{kj}$ can be interpreted as the log-fold difference in $\mathbb{E}W_{\cdot j}$ across groups that differ by one unit in $X_{\cdot k}$, minus the corresponding log-fold difference in $\mathbb{E}W_{\cdot j^{\prime}}$. That is, $\beta_{kj}$ can only be interpreted relative to $\beta_{kj^{\prime}}$.
@@
Identifiability can also be imposed relative to all categories, rather than relative to a specific category. We first consider constraining $\text{median}_{j=1,\ldots,J}(\beta_{kj})=0$ for all $k$. The robustness of the median is attractive, as no single large effect size will impact the median (Supplementary Figure 1). However, the median is not a differentiable function of $\beta_{k}$ (its total derivative is not guaranteed to exist), which substantially complicates inference (see Section 5). Accordingly, we propose a smooth approximation to the median that shares its robustness properties: the pseudo-Huber loss $g_{p}(\beta_{k})=\arg\min_{c}\sum_{j=1}^{J}h_{\gamma}\left(\beta_{kj}-c\right)$ for $h_{\gamma}(x)=\gamma^{2}\left(\sqrt{1+\left(x/\gamma\right)^{2}}-1\right)$ for prespecified parameter $\gamma>0$ (Barron, 2019). The pseudo-Huber loss is approximately quadratic in a neighborhood of 0 and asymptotically approaches the function $|x|/\gamma$ for large $x$. The parameter $\gamma$ controls the size of the quadratic region around zero: as $\gamma\downarrow 0$, $g_{p}(\beta_{k})$ approaches the median, and as $\gamma\rightarrow\infty$, it approaches the mean. We propose $\gamma=0.1$ as a reasonable default for many applications, though empirically we find that care is warranted when covariates are on very different scales. The following example illustrates that enforcing identifiability via the smoothed median constraint does not imply any structural assumptions on $\beta$.
Example 2: Similar to Example 1, if $g_{p}(\beta_{k})$ is known to equal zero (e.g., $\beta_{k}$ is known to be sparse or symmetric), then $\beta_{kj}$ can be interpreted as the log-fold difference in $\mathbb{E}W_{\cdot j}$ across groups that differ by one unit in $X_{\cdot k}$. Without this assumption, applying the constraint $g_{p}(\beta_{k})=0$ means that $\beta_{kj}$ should be interpreted as the log-fold difference in $\mathbb{E}W_{\cdot j}$ across groups that differ by one unit in $X_{\cdot k}$ relative to the typical such log-fold difference across taxa.
Further illustration of the advantages of estimating fold-differences relative to smoothed medians are given in Supplementary Figure 1. We propose $g_{p}$ as the default constraint function and give a detailed example of interpreting partially identifiable estimands without structural assumptions on $\beta$ in Section 7. That said, any smooth $g:\mathbb{R}^{J}\rightarrow\mathbb{R}$ satisfying $g(\beta_{k}+\alpha)=g(\beta_{k})+\alpha$ is possible, and the most appropriate constraint will depend on the scientific setting.